# Lrq-Fact: LLM-Generated Relevant Questions for Multimodal Fact-Checking

### v1

@@ A.4 Instruct Prompt for Lrq-Fact
Image-Focused Questions Prompt: Given the following news article [news text], generate up to [number of questions] questions that are relevant to the news article and focused on elements visible in the image. These questions should help clarify the image content, assess its authenticity, and explore its relation to the news article.
Instructions for Question Generation:
Focus on generating high-quality questions that are directly relevant to both the news article and the visual elements in the image. The questions should investigate visible interactions, settings, actions, any text or symbols, and specific individuals or objects mentioned in the article. Additionally, include questions that assess the authenticity of the image, such as whether the image could have been generated by AI, or whether there are any unusual or suspicious elements in the image.
Aim to Generate:
- The most important questions that will help determine whether the image supports or contradicts the context provided by the news article.
- Questions that may uncover discrepancies or inconsistencies between the image and the news article.
- Questions that assess the authenticity or realism of the image, particularly in relation to the article’s content.
Avoid asking for:
- Personal details not visible in the image.
- Information that cannot be inferred from the image itself.
@@ 5. What emotions does this image evoke?
Figure 9: Structured prompt to generate relevant image-focused questions.
Image-Focused Answers Prompt: You are an advanced AI model with access to a vast repository of knowledge and the capability to analyze both text [news text] and image [image]. Ensure your answer is directly related to the question [generated questions] and it is based on the image. Answers: 1. , 2. , …
Figure 10: Structured prompt to generate answers for the image-focused questions.
Text-Focused Questions Prompt: Given the following news article [news text], analyze the text and formulate up to [number of questions] questions that probe the accuracy and verifiability of the information contained in the article. These questions should be designed to identify potential inaccuracies or areas that can be confirmed or challenged based on general knowledge or the text itself.
Instructions for Question Generation:
@@
Figure 11: Structured prompt to generate relevant text-focused questions.
Text-Focused Answers Prompt: You are an advanced AI model with access to a vast repository of knowledge. Your task is to answer the following questions [generated questions] based on your built-in knowledge. While a news article [news text] is provided for context, you must answer the questions solely based on your own knowledge and not refer to the article’s content.
Instructions for Answering:
@@
Figure 12: Structured prompt to generate answers for the relevant text-focused questions.
@@ 2. Image Description: [image description]
3. Generated Image-Focused Questions and Answers: [generated image-focused QAs]
@@
4. Generated Text-Focused Questions and Answers: [generated text-focused QAs]
@@ 8. Provide a Detailed Explanation: …
1. Final Judgment: Fake
@@ 2. Label: Visual Veracity Distortion
1. Final Judgment: [Real or Fake]
2. Label: [Select one: Textual Veracity Distortion, Visual Veracity Distortion, Mismatch, Real]
3. Explanation: [Provide your explanation here]

### v2

@@ A.5 Dataset Descriptions and Details
Visual Questions Prompt: Given the following news article [news text], generate up to [number of questions] questions that are directly based on the news article and are designed to explore visual elements that could be present in an image related to the article. Instructions for Question Generation:
Focus on generating questions that are directly relevant to the news article and the visual elements that could be present in an image. The questions should examine visible interactions, settings, actions, text, symbols, and specific objects mentioned in the article. Additionally, include questions that assess the authenticity of the image, such as whether it could have been AI-generated or contains any unusual or suspicious elements.
Avoid the following in your Questions:
- Do not mention any names.
- Do not ask questions about identification .
- Do not ask about personal details.
- Do not ask compound questions in a single sentence.
@@ 5. What emotions does this image evoke?
6. What are the main objects or elements visible in this image?
7. What unusual elements in the image might suggest digital manipulation or artificial creation?
@@
Figure 12: Structured prompt to generate relevant visual questions.
Visual Answers Prompt: You are an advanced AI model with access to a vast repository of knowledge and the capability of answering image questions. Your task is to answer the following questions [generated questions] based on the image [image]. While a news article [news text] is provided for context, you must answer the questions solely based on the image and not refer to the article’s content. Instructions for Answer Generation:
- Provide accurate, clear, and concise answers to each question.
- Your responses should be based entirely on the image.
- Do not reference or rely on the content of the provided news article when forming your answers.
- Each answer should be directly relevant to the question asked.
Avoid the following in your Answers:
- Provide accurate, clear, and concise answers to each question.
- Your responses should be based entirely on the image.
- Do not reference or rely on the content of the provided news article when forming your answers.
- Each answer should be directly relevant to the question asked.
Answers: 1. , 2. , …
Figure 13: Structured prompt to generate answers for the visual questions.
Textual Questions Prompt: Given the following news article [news text], analyze the text and formulate up to [number of questions] questions that probe the accuracy and verifiability of the information contained in the article. These questions should be designed to identify potential inaccuracies or areas that can be confirmed or challenged based on general knowledge or the text itself. Instructions for Question Generation:
@@
Figure 14: Structured prompt to generate relevant textual questions.
Textual Answers (w/o Evidence) Prompt: You are an advanced AI model with access to a vast repository of knowledge. Your task is to answer the following questions [generated questions] based on your built-in knowledge. While a news article [news text] is provided for context, you must answer the questions solely based on your own knowledge and not refer to the article’s content. Instructions for Answering:
@@
Figure 15: Structured prompt to generate answers based on llm-knowledge for the relevant textual questions.
Textual Answers (w/Evidence) Prompt: You are an advanced AI tasked with evaluating the authenticity of a news article. Your task is to answer the following questions [generated questions] based on the provided factual document [evidence]. While a news article [news text] is provided for context, you must answer the questions solely based on the provided factual document and not refer to the article’s content. Instructions for Answering:
Provide accurate, clear, and concise answers to each question. Your responses should be based entirely on the provided factual document. If there was no factual answer for the question use your built-in knowledge to answer the question. Do not reference or rely on the content of the provided news article when forming your answers. Each answer should be directly relevant to the question asked.
Answers: 1. , 2. , …
Figure 16: Structured prompt to generate answers based on factual evidence for the relevant textual questions.
Questions Quality Prompt: You are an advanced AI tasked with evaluating the authenticity of a news article and its accompanying image. Your objective is to determine whether the provided questions [generated questions] effectively assess the accuracy, credibility, and reliability of the news article text. Instructions:
Assess each question based on the following expert fact-checking criteria:
1. Critical Thinking and Skepticism: Does the question challenge assumptions, probe deeper into claims, and avoid taking information at face value?
2. Analytical Depth: Does it break down complex statements into verifiable components?
3. Systematic Approach: Does it follow a structured methodology in assessing sources, claims, and evidence?
4. Precision & Specificity: Is it clear, direct, and free from vague or overly broad wording?
5. Factual Accuracy: Does it focus on verifying evidence, checking primary sources, and detecting misinformation?
6. Logical Consistency: Does it help identify contradictions, misleading narratives, or inconsistencies?
7. Source Credibility & Bias Detection: Does it evaluate the reliability of cited sources and potential biases?
8. Context Awareness: Does it consider the broader context surrounding the claim?
9. Comparative Thinking: Does it encourage cross-referencing with established facts or alternative perspectives?
10. Repeatability & Objectivity: Can the question be applied consistently across different cases without personal bias?
Rating Scale:
- Relevant: The question is precise, well-structured, and effectively assesses factual accuracy, credibility, and logical consistency.
- Irrelevant: The question is vague, lacks depth, or fails to critically probe the credibility and factuality.
Answers:
Q1: [Relevant or Irrelevant]
Q2: [Relevant or Irrelevant]
Figure 17: Structured prompt to evaluate the quality of generated questions.
@@ 2. Image Description: [image description]
3. Generated Visual Questions and Answers: [generated visual FCQs]
@@
4. Generated Textual Questions and Answers: [generated textual FCQs]
@@ 8. Provide a Detailed Explanation: …
1. Final Judgment:
Fake
@@ 2. Label: Visual Veracity Distortion
1. Final Judgment: [Real or Fake] 2. Label: [Select one: Textual Veracity Distortion, Visual Veracity Distortion, Mismatch, Real] 3. Explanation: [Provide your explanation here]
@@
Appendix B Acknowledgment of AI Assistance in Writing and Revision
We utilized ChatGPT-4 for revising and enhancing sections of this paper.